earth, which is an electrical conductor, carries a net charge of -4.3 x 10⁵ distributed approximately uniformly over it surface. The electric field at Earth's surface is -8.43x10¹⁰ C/m². find the electric field in the earth surface. what is the meaning of the minus signin your final answer?


1
Expert's answer
2021-05-05T14:08:41-0400

"E=k\frac{q}{r^2}=9\cdot10^9\cdot \frac{-4.3\cdot10^5}{6400000^2}=-94.5\ (V/m)"


The vector of the electric field strength is directed to the Earth's surface.